# Polynomial stability of a coupled wave-heat network

### v1

@@ 1. Introduction
$E(t)=\frac{1}{2}\sum_{k=1}^{3}\int_{0}^{1}|y^{k}_{x}(x,t)|^{2}+|y^{k}_{t}(x,t)|^{2}\,\mathrm{d}x+\frac{1}{2}\sum_{k=1}^{2}\int_{0}^{1}|w^{k}(x,t)|^{2}\,\mathrm{d}x$
@@ 2. Boundary nodes and an abstract stability result
The following is a special case of [24, Thm. 3.1]. The result makes it possible to deduce a growth bound for the resolvent of $A$ from information about the boundary nodes $(G_{1}+K_{1},L_{1},K_{1})$ and $(G_{2},L_{2},K_{2})$. Here we may think of $(G_{1}+K_{1},L_{1},K_{1})$ as a boundary-damped version of the boundary node $(G_{1},L_{1},K_{1})$. Note that in the coupled system this boundary-damping is replaced by the interconnection with the boundary node $(G_{2},L_{2},K_{2})$.
@@
Note that [24, Thm. 3.1] is in fact more general than this. In particular, in [24, Thm. 3.1] the space $U$ is allowed to be different for each of two boundary nodes, and in addition we may replace (2.4) by the weaker condition that the norms of the two resolvents can be estimated from above along the imaginary axis by functions which need not be bounded.
@@ 3. Analysis of the wave-heat network
We begin by expressing the dynamics of the our wave-heart network as an abstract Cauchy problem on a Hilbert space $X=L^{2}(0,1)^{7}$. To this end, we define
@@
$Ax=(v_{1}^{\prime},u_{1}^{\prime},v_{2}^{\prime},u_{2}^{\prime},w_{1}^{\prime\prime},w_{2}^{\prime\prime},w_{3}^{\prime\prime})^{\top}$
@@
$L_{2}\begin{pmatrix}w_{1}\\ w_{2}\\ w_{3}\end{pmatrix}=\begin{pmatrix}w_{1}^{\prime\prime}\\ w_{2}^{\prime\prime}\\ w_{3}^{\prime\prime}\end{pmatrix}$
@@
We now show that $(G_{1},L_{1},K_{1})$ and $(G_{2},L_{2},K_{2})$ are impedance passive boundary nodes, and that the restrictions $A_{0}$ and $A_{2}$ of $L_{1}$ to $\operatorname*{Ker}(G_{1}+K_{1})$ and of $L_{2}$ to $\operatorname*{Ker}G_{2}$, respectively, have uniformly bounded resolvents along the imaginary axis. Furthermore, we derive bounds for the transfer function $P_{2}$ of the second node. Once these objectives have been met, Theorem 1.1 will be a straightforward consequence of Theorem 2.1.
@@ 3.1. Analysis of the wave network
In this section we analyse the wave network and show that $(G_{1},L_{1},K_{1})$ defines an impedance passive boundary node on $(\mathbb{C}^{3},X_{1},\mathbb{C}^{3})$, where $X_{1}=L^{2}(0,1)^{4}$. From the above definitions it is clear that $G_{1},K_{1}\in\mathcal{B}(D(L_{1}),\mathbb{C}^{3})$, and since $G_{1}$ is surjective and $\mathbb{C}^{3}$ is finite-dimensional we know that $G_{1}$ has a bounded right-inverse. We note that the restriction $A_{1}$ of $L_{1}$ to $\operatorname*{Ker}G_{1}$ has the structure $A_{1}=\operatorname*{diag}(A_{1,1},A_{1,2})$, where $A_{1,1},A_{1,2}\colon(u,v)^{\top}\mapsto(v^{\prime},u^{\prime})^{\top}$ with respective domains $D(A_{1,1})=\{(u,v)^{\top}\in H^{1}(0,1)^{2}:u(1)=0,v(0)=0\}$ and $D(A_{1,2})=H^{1}(0,1)\times H_{0}^{1}(0,1)$. Since both $A_{1,1}$ and $A_{1,2}$ generate strongly continuous contraction semigroups on $L^{2}(0,1)^{2}$, $A_{1}$ generates a contraction semigroup on $X_{1}$. Thus $(G_{1},L_{1},K_{1})$ is a boundary node. A straightforward computation using integration by parts shows that $\operatorname*{Re}\langle L_{1}x,x\rangle=\operatorname*{Re}\langle G_{1}x,K_{1}x\rangle$ for all $x\in D(L_{1})$, so the boundary node $(G_{1},L_{1},K_{1})$ is impedance passive. Arguing as for instance in [17, Ex. 9.2.1] for the wave equations along each of the blue edges in Figure 2 it is possible to show that the semigroup generated by the restriction $A_{0}$ of $L_{1}$ to $\operatorname*{Ker}(G_{1}+K_{1})$ is uniformly exponentially stable. In fact, by adapting the argument given in [9, Sec. 1] we see that this semigroup is even nilpotent in our case. Either way, it follows that $i\mathbb{R}\subseteq\rho(A_{0})$ and $\sup_{s\in\mathbb{R}}\|(is-A_{0})^{-1}\|<\infty$.
@@ 3.2. Analysis of the heat network
$q_{1}^{k}(s)^{2}-q_{2}^{k}(s)^{2}=\frac{\beta_{k}|s|^{2}}{2}\frac{\sinh a_{k}(s)^{2}-\sin a_{k}(s)^{2}}{\sinh a_{k}(s)^{2}+\sin a_{k}(s)^{2}},$
@@ 4. Energy decay
In this final section we combine the results presented in Section 2 in order to prove our main result.
@@
We consider the abstract Cauchy problem corresponding to our wave-heat system as explained in Section 3. It follows from the observations made in Section 3.1 and from Proposition 3.2 that the operator $A$ defined in Section 3 generates a contraction semigroup $(T(t))_{t\geq 0}$ on the space $X=X_{1}\times X_{2}$. Proposition 3.2 also tells us that $\|P_{2}(1+is)\|^{2}\lesssim 1+|s|$ for $s\in\mathbb{R}$ and that, given any $\varepsilon>0$, there exists $c_{\varepsilon}$ such that $\operatorname*{Re}P_{2}(is)\geq c_{\varepsilon}(1+|s|^{1/2})I$ for all $s\in\mathbb{R}\setminus(-\varepsilon,\varepsilon)$. It follows from Theorem 2.1 that $\{is:s\in\mathbb{R},|s|>\varepsilon\}\subseteq\rho(A)$ and $\|(is-A)^{-1}\|\lesssim 1+|s|^{1/2}$ for $|s|>\varepsilon$, for any given $\varepsilon>0$. Next we wish to establish that $0\in\rho(A)$. We begin by computing $P_{0}(0)$, where $P_{0}$ is the transfer function of the boundary node $(G_{1}+K_{1},L_{1},K_{1})$. To this end we note that if $u=(a,b,c)^{\top}\in\mathbb{C}^{3}$ then $P_{0}(0)=(-v_{1}(1),-v_{2}(0),-v_{2}(1))^{\top}$, where $u_{1},v_{1},u_{2},v_{2}\in H^{1}(0,1)$ satisfy $u_{1}^{\prime}=v_{1}^{\prime}=u_{2}^{\prime}=v_{2}^{\prime}=0$ as well as $v_{1}(0)=0$, $-u_{1}(1)-v_{1}(1)=a$, $u_{2}(0)-v_{2}(0)=b$ and $-u_{2}(1)-v_{2}(1)=c$. Thus $u_{1},v_{1},u_{2},v_{2}$ are all constant, $v_{1}=0$, and $v_{2}=-(b+c)/2$. It follows that
$P_{0}(0)=\begin{pmatrix}0&0&0\\ 0&1/2&1/2\\ 0&1/2&1/2\end{pmatrix}.$
From the proof of Proposition 3.2 we know that
$P_{2}(0)=\begin{pmatrix}\beta_{1}+\beta_{3}&-\beta_{1}&-\beta_{3}\\ -\beta_{1}&\beta_{1}+\beta_{2}&-\beta_{2}\\ -\beta_{3}&-\beta_{2}&\beta_{2}+\beta_{3}\end{pmatrix},$
and hence
$I+P_{2}(0)P_{0}(0)=\frac{1}{2}\begin{pmatrix}2&-\beta_{1}-\beta_{3}&-\beta_{1}-\beta_{3}\\ 0&\beta_{1}+2&\beta_{1}\\ 0&\beta_{3}&\beta_{3}+2\end{pmatrix}.$
In particular, the matrix $I+P_{2}(0)P_{0}(0)$ is non-singular. It follows from [24, Rem. 3.10] that $0\in\rho(A)$. By choosing $\varepsilon>0$ in the previous part of the argument to be sufficiently small we may therefore deduce that $i\mathbb{R}\subseteq\rho(A)$ with $\|(is-A)^{-1}\|\lesssim 1+|s|^{1/2}$ for $s\in\mathbb{R}$. Thus the semigroup $(T(t))_{t\geq 0}$ is strongly stable by an application of the theorem of Arendt, Batty, Lyubich and Vũ [14, Thm. V.2.21], and by [8, Thm. 2.4] we obtain $\|T(t)z_{0}\|=o(t^{-2})$ as $t\to\infty$ for all $z_{0}\in D(A)$. Since the energy of a solution with initial data $z_{0}\in X$ is given by $E(t)=\frac{1}{2}\|T(t)z_{0}\|^{2}$ for all $t\geq 0$, the conclusions of Theorem 1.1 follow immediately. ∎
@@
We may alternatively consider the wave-heat system with the Neumann boundary condition $y_{x}^{1}(t,0)=0$ for $t>0$ at the left end of the first wave equation in Figure 1. The analysis is largely the same as in our case, with the important difference that the point 00 becomes a spectral point of the operator $A$.

### v2

@@ 1. Introduction
$E(t)=\frac{1}{2}\sum_{k=1}^{2}\int_{0}^{1}|y^{k}_{x}(x,t)|^{2}+|y^{k}_{t}(x,t)|^{2}\,\mathrm{d}x+\frac{1}{2}\sum_{k=1}^{3}\int_{0}^{1}|w^{k}(x,t)|^{2}\,\mathrm{d}x$
@@ 2. Boundary nodes and an abstract stability result
The following is a special case of [24, Thm. 3.1]. The result makes it possible to deduce a growth bound for the resolvent of $A$ from information about the boundary nodes $(G_{1}+K_{1},L_{1},K_{1})$ and $(G_{2},L_{2},K_{2})$. Here we may think of $(G_{1}+K_{1},L_{1},K_{1})$ as a boundary-damped version of the boundary node $(G_{1},L_{1},K_{1})$. For instance, in our wave-heat network $(G_{1}+K_{1},L_{1},K_{1})$ describes a wave network with dissipative boundary conditions. Note that in the coupled system this boundary-damping is replaced by the interconnection with the boundary node $(G_{2},L_{2},K_{2})$.
@@
Note that [24, Thm. 3.1] is in fact more general than this. In particular, in [24, Thm. 3.1] the space $U$ is allowed to be different for each of the two boundary nodes, and in addition we may replace (2.3) by the weaker condition that the norms of the two resolvents can be estimated from above along the imaginary axis by functions which need not be bounded.
@@ 3. Analysis of the wave-heat network
We begin by expressing the dynamics of the wave-heat network as an abstract Cauchy problem on a Hilbert space $X=L^{2}(0,1)^{7}$. To this end, we define
@@
$Ax=(v_{1}^{\prime},u_{1}^{\prime},v_{2}^{\prime},u_{2}^{\prime},\beta_{1}w_{1}^{\prime\prime},\beta_{2}w_{2}^{\prime\prime},\beta_{3}w_{3}^{\prime\prime})^{\top}$
@@
$L_{2}\begin{pmatrix}w_{1}\\ w_{2}\\ w_{3}\end{pmatrix}=\begin{pmatrix}\beta_{1}w_{1}^{\prime\prime}\\ \beta_{2}w_{2}^{\prime\prime}\\ \beta_{3}w_{3}^{\prime\prime}\end{pmatrix}$
@@
We now show that $(G_{1},L_{1},K_{1})$ and $(G_{2},L_{2},K_{2})$ are impedance passive boundary nodes, and that the restrictions $A_{0}$ and $A_{2}$ of $L_{1}$ to $\operatorname*{Ker}(G_{1}+K_{1})$ and of $L_{2}$ to $\operatorname*{Ker}G_{2}$, respectively, have uniformly bounded resolvents along the imaginary axis. Furthermore, we derive bounds for the transfer function $P_{2}$ of the second node. Once these objectives have been met, Theorem 1.1 will be a straightforward consequence of Theorem 2.1 and the stability theory of strongly continuous semigroups.
@@ 3.1. Analysis of the wave network
In this section we analyse the wave network and show that $(G_{1},L_{1},K_{1})$ defines an impedance passive boundary node on $(\mathbb{C}^{3},X_{1},\mathbb{C}^{3})$, where $X_{1}=L^{2}(0,1)^{4}$. From the above definitions it is clear that $G_{1},K_{1}\in\mathcal{B}(D(L_{1}),\mathbb{C}^{3})$, and since $G_{1}$ is surjective and $\mathbb{C}^{3}$ is finite-dimensional we know that $G_{1}$ has a bounded right-inverse. We note that the restriction $A_{1}$ of $L_{1}$ to $\operatorname*{Ker}G_{1}$ has the structure $A_{1}=\operatorname*{diag}(A_{1,1},A_{1,2})$, where $A_{1,1},A_{1,2}\colon(u,v)^{\top}\mapsto(v^{\prime},u^{\prime})^{\top}$ with respective domains $D(A_{1,1})=\{(u,v)^{\top}\in H^{1}(0,1)^{2}:u(1)=0,v(0)=0\}$ and $D(A_{1,2})=H_{0}^{1}(0,1)\times H^{1}(0,1)$. Since both $A_{1,1}$ and $A_{1,2}$ generate strongly continuous contraction semigroups on $L^{2}(0,1)^{2}$, $A_{1}$ generates a contraction semigroup on $X_{1}$. Thus $(G_{1},L_{1},K_{1})$ is a boundary node. A straightforward computation using integration by parts shows that $\operatorname*{Re}\langle L_{1}x,x\rangle=\operatorname*{Re}\langle G_{1}x,K_{1}x\rangle$ for all $x\in D(L_{1})$, so the boundary node $(G_{1},L_{1},K_{1})$ is impedance passive. Arguing as for instance in [17, Ex. 9.2.1] for the wave equations along each of the blue edges in Figure 2 it is possible to show that the semigroup generated by the restriction $A_{0}$ of $L_{1}$ to $\operatorname*{Ker}(G_{1}+K_{1})$ is uniformly exponentially stable. In fact, by adapting the argument given in [9, Sec. 1] we see that this semigroup is even nilpotent in our case. Either way, it follows that $i\mathbb{R}\subseteq\rho(A_{0})$ and $\sup_{s\in\mathbb{R}}\|(is-A_{0})^{-1}\|<\infty$.
@@ 3.2. Analysis of the heat network
$q_{1}^{k}(s)^{2}-q_{2}^{k}(s)^{2}=\frac{\beta_{k}|s|}{2}\frac{\sinh^{2}a_{k}(s)-\sin^{2}a_{k}(s)}{\sinh^{2}a_{k}(s)+\sin^{2}a_{k}(s)},$
@@ 4. Energy decay
In this final section we combine the results presented in Sections 2 and 3 in order to prove our main result.
@@
We consider the abstract Cauchy problem corresponding to our wave-heat system as explained in Section 3. The observations made in Section 3.1 together with Proposition 3.2 imply that the boundary nodes $(G_{1},L_{1},K_{1})$ and $(G_{2},L_{2},K_{2})$ satisfy the conditions of Theorem 2.1 with $E=\mathbb{R}\setminus(-\varepsilon,\varepsilon)$ for any $\varepsilon>0$. Therefore the operator $A$ defined in Section 3 generates a contraction semigroup $(T(t))_{t\geq 0}$ on the space $X=X_{1}\times X_{2}$. Proposition 3.2 also tells us that $\|P_{2}(1+is)\|^{2}\lesssim 1+|s|$ for $s\in\mathbb{R}$ and that, given any $\varepsilon>0$, there exists $c_{\varepsilon}$ such that $\operatorname*{Re}P_{2}(is)\geq c_{\varepsilon}(1+|s|^{1/2})I$ for all $s\in\mathbb{R}\setminus(-\varepsilon,\varepsilon)$. It follows from Theorem 2.1 that $\{is:s\in\mathbb{R},|s|>\varepsilon\}\subseteq\rho(A)$ and $\|(is-A)^{-1}\|\lesssim 1+|s|^{1/2}$ for $|s|>\varepsilon$, for any given $\varepsilon>0$. Next we establish that $0\in\rho(A)$. We first note that since both $A_{1}$ and $A_{2}$ have compact resolvent, it follows from [24, Thm. 3.7] that $A$, too, has compact resolvent. It therefore suffices to show that $A$ is injective. To this end, suppose that $x=(u_{1},v_{1},u_{2},v_{2},w_{1},w_{2},w_{3})^{\top}\in\operatorname*{Ker}A$. Then the functions $u_{k},v_{k}$ are constant for $k=1,2$, and the functions $w_{k}$ are affine for $k=1,2,3$. Moreover, the coupling conditions in $D(A)$ give
$v_{1}(0)=0,\quad$ $v_{1}(1)=w_{1}(0)=w_{3}(1),\quad u_{1}(1)-\beta_{1}w_{1}^{\prime}(0)+\beta_{3}w_{3}^{\prime}(1)=0,$
$w_{1}(1)=w_{2}(0)=v_{2}(0),\quad\beta_{1}w_{1}^{\prime}(1)-\beta_{2}w_{2}^{\prime}(0)-u_{2}(0)=0,$
$w_{2}(1)=w_{3}(0)=v_{2}(1),\quad\beta_{2}w_{2}^{\prime}(1)-\beta_{3}w_{3}^{\prime}(0)+u_{2}(1)=0.$
Since $\beta_{1},\beta_{2},\beta_{3}>0$, it follows easily that $x=0$. Hence $A$ is injective, and therefore invertible. By choosing $\varepsilon>0$ in the previous part of the argument to be sufficiently small we may therefore deduce that $i\mathbb{R}\subseteq\rho(A)$ with $\|(is-A)^{-1}\|\lesssim 1+|s|^{1/2}$ for $s\in\mathbb{R}$. Thus the semigroup $(T(t))_{t\geq 0}$ is strongly stable by an application of the theorem of Arendt, Batty, Lyubich and Vũ [14, Thm. V.2.21], and by [8, Thm. 2.4] we obtain $\|T(t)z_{0}\|=o(t^{-2})$ as $t\to\infty$ for all $z_{0}\in D(A)$. Since the energy of a solution with initial data $z_{0}\in X$ is given by $E(t)=\frac{1}{2}\|T(t)z_{0}\|^{2}$ for all $t\geq 0$, the conclusions of Theorem 1.1 follow immediately. ∎
@@
We may alternatively consider the wave-heat system with the Neumann boundary condition $y_{x}^{1}(t,0)=0$ for $t>0$ at the left end of the first wave equation in Figure 1. The analysis is largely the same as in our case, with the important difference that the point 00 becomes an eigenvalue of the operator $A$.